NathanielFreel:

Marneshia walked 3/8 of a mile in 3/5 of an hour. What equation can be used to calculate her unit rate in miles per hour?

NathanielFreel:

If anyone answers this, I will reward them greatly!

lowkey:

How do you think you start the equation?

NathanielFreel:

you start by multipling 3/8 by 3/5 in a cross, called cross multipling.

NathanielFreel:

Right? I am not good in math.

lowkey:

The way I did it is, I divide 3/8 and 3/5 together. This way we can find the unit rate by every mile

NathanielFreel:

OHHH!!!! Ok... I see it! You DIVIDED! my bad! You see... I am in home schooling, and I am in virtual school, so it is hard for me to do math!

lowkey:

Yeah! and do you know the main rule when dividing fractions? 3/8 divided by 3/5=

NathanielFreel:

Fractions is basicly dividing both numbers on top and bottom... so 3/8 would be 2.6 and 3/5 would be 1.6 and the desimals would be divided and would equal 1.625!

lowkey:

No, sorry it doesn't work that way! Don't do it like that okay x'd Usually it's like this so for example, \[\frac{ 1 }{ 2 }/\frac{ 3 }{ 4 }= \frac{ 1 }{ 2 }*\frac{ 4 }{ 3 }=\frac{ 4 }{ 5 }\] When you divide fractions, you have to multiply the first fraction by the reciprocal of the second function. The reciprocal is the opposite number, which is flipped, so that is why 3/4 in this situation went to 4/3.

lowkey:

For this situation, you would multiple 3/8 and the reciprocal of 3/5, which is 5/3, and multiply it together. The final answer would be 5/8.

justjm:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @NathanielFreel Fractions is basicly dividing both numbers on top and bottom... so 3/8 would be 2.6 and 3/5 would be 1.6 and the desimals would be divided and would equal 1.625! \(\color{#0cbb34}{\text{End of Quote}}\) Oh dear XD